# Locally-Supervised Global Image Restoration

### v1

@@ 2.1 Problem formulation
The optimal restoration function $\hat{\Phi}\colon\mathds{R}^{I}\to\mathds{R}^{I}$ is defined as the minimizer of the supervised loss function
$\mathcal{L}(\Phi)\coloneqq\mathbb{E}[\lVert\Phi(X_{\Omega})-X\rVert^{2}]\,$, (2.1)
where the minimum is taken over all measurable functions $\Phi\colon\mathds{R}^{I}\to\mathds{R}^{I}$. It is well known that the minimizer of (2.1) is given by the conditional expectation $\hat{\Phi}=\mathbb{E}[X|X_{\Omega}]$.
Computing the exact minimizer in (2.1) requires exact knowledge of the joint distribution $\pi$, which is usually not available. In the supervised learning paradigm, the minimizer of (2.1) is instead approximated by minimizing the empirical loss $\sum_{n=1}^{N}\lVert\Phi(y_{n})-x_{n}\rVert^{2}$ using samples $(y_{n},x_{n})$ drawn independently from the joint distribution. This, however, requires fully sampled ground truth images $x_{n}$, which again might not be available in practice.
@@ 2.2 Self-supervised upsampling
In the context of image upsampling or inpainting, the self-supervision paradigm selects a second subsampling set $\Gamma\subseteq\Omega$, which further degrades the partially observed data. This procedure defines synthetic training pairs $(P_{\Lambda}y_{n},y_{n})$, where the original partially observed image $y_{n}$ serves as the ground truth, and $P_{\Lambda}y_{n}$ acts as the input to the model. In the context of image restoration with known masks, the main theoretical justification is that $\mathbb{E}[X|M_{\Omega}\odot X]$ can in fact be constructed from data $(P_{\Lambda}y,y)$. For convenience of the reader, we state a main theoretical result in this context due to [9], where multiplicative variants of noisier2noise and their relation to SSDU (self-supervised learning via data undersampling, proposed in [16]) have been studied.
Proposition 2.1 (Recovery guarantee for SSDU).
Let $\Omega,\Lambda\subseteq I$ be random subsets with $\mathbb{E}[M_{\Omega}]>0$ and $\mathbb{E}[M_{\Lambda}]<1$. Then
$M_{I\setminus(\Omega\cap\Lambda)}\odot\mathbb{E}[X|M_{\Lambda}\odot X_{\Omega}]\\ =M_{I\setminus(\Omega\cap\Lambda)}\odot\Bigl(\operatorname*{arg\,min}_{\Phi}\mathbb{E}\bigl[\lVert M_{I\setminus\Lambda}\odot\bigl(M_{\Omega}\odot\Phi(M_{\Lambda}\odot X_{\Omega})-X_{\Omega}\bigr)\rVert^{2}|M_{\Lambda}\odot X_{\Omega}\bigr]\Bigr)\,.$ (2.2)
Proposition 2.1 states that on the pixels outside $\Omega\cap\Lambda$ the ideal estimator $\mathbb{E}[X|M_{\Lambda}\odot X_{\Omega}]$ for given further masked data $M_{\Lambda}\odot X_{\Omega}$, can be found using only access to $X_{\Omega}$ instead of $X$ during training. In our opinion, this is quite remarkable. However, a key assumption is the randomness of the set $\Omega$ that in expectation covers the whole region $I$. In many practical scenarios, however, the set $\Omega$ may be deterministic and fixed, and moreover does not cover the entire imaging domain. Our method can be seen as an extension to such a non-random, incomplete scenario. In particular, it allows for a fixed, non-random design. To obtain sufficient information, we exploit invariances in the distribution of $X$.
@@ 3.2 Main results
The invariance of $\mathbb{E}[X|X_{\Omega}]$ follows from Lemma 3.4. Further, from the $\mathcal{T}$-invariance of $\mathbb{E}[X|X_{\Omega}]$ and conditions (A3), (A4), we get
@@ 4.2 Sparse-Dense OR-PAM
Other invariances are also possible, for example, rotating the image by multiples of $90$ degrees; however, in this paper, we stick to translation invariances only.
@@ 4.3 Implementation details
To ensure that the network satisfies the condition of a $\mathcal{T}$-equivariant function, we analyze the translation invariance properties of the architecture in detail:
  • Convolutional layers are translation-equivariant because the same filter is applied with shared weights across all spatial positions, so a shift in the input produces an equivalent shift in the feature map. With standard settings, this property holds only in the interior of the image; at the boundaries, padding and edge effects may break perfect invariance. Therefore, all convolutional layers use padding_mode=circular, which ensures that the convolution operator is translation-equivariant with periodic boundary conditions. For convolutions with stride $s=1$, translation invariance holds for all integer shifts. However, the encoder path contains three strided convolutional layers (stride=2) acting as downsampling operators. Similarly, the decoder path contains three strided transposed convolutional layers (stride=2) acting as upsampling operators. For these layers, translation invariance is guaranteed only for translations that are multiples of the stride. Otherwise, aliasing effects occur, which break exact invariance. Since there are three downsampling operations in the encoder, the overall stride is $2\cdot 2\cdot 2=8$.
@@ 4.4 Numerical Studies
Figure 4.7: Visualization of Superior performance of sparse-dense sampling over patch-wise learning. From left to right: ground truth (image 1); downsampled image (image 2); upsampling with a patch-based approach (image 3); upsampling with sparse-dense sampling (image 4).
@@ 5 Conclusion
Future work will focus on studying the influence of noise and exploring other invariances, such as rotation and mirroring. We will also investigate the extent to which the sampling rate can be reduced. Finally, we plan to extend our method to random sampling and self-supervision with noisy data, where not all samples are available.

### v2

@@ 1 Introduction
More precise formulations of the required elements are given in Condition 3.9. Note these assumptions, in particular, require that $B\cap\Omega^{c}\neq\emptyset$. Otherwise we would have $B\subseteq\Omega$, in which case no supervised data for the prediction task are collected at all. We also point out that $\Omega$ and $B$ have very different roles. On $\Omega$ we collect observed data for the actual inverse problem under study. On the other hand, on $B$ we collect additional data only during the training phase. Without essential changes, we could assume $B\cap\Omega=\emptyset$, as the supervision pixels in $B\cap\Omega$ do not provide additional information. However, we prefer the present setting where we formally aim at estimating $X$ given $X_{\Omega}$ instead of estimating $X_{\Omega^{c}}$ given $X_{\Omega}$, which in the noise-free inpainting case is equivalent. In particular, this excludes the extreme case $B\subseteq\Omega$, and $B\cap\Omega^{c}$ is required to have at least one element. However, as $(T_{\ell}(B))_{\ell=1}^{L}$ is assumed to form a partition of $I$, the set of invariances has to be large enough. Because $(T_{\ell}(B\setminus\Omega))_{\ell=1}^{L}$ forms a partition of $I\setminus\Omega$, this gives the requirement $L\lvert B\setminus\Omega\rvert=\lvert I\setminus\Omega\rvert$.
@@
From a practical point of view, our strategy requires more data pairs $(Y,X_{B})$ and thus additional measurements $X_{B}$ on the supervision region $B$ during training compared to purely self-supervised approaches. However, the set of additional measurements $B\setminus\Omega$ can be much smaller than the full supervision $I\setminus\Omega$ needed in fully supervised approaches based on $(Y,X)$. Depending on the available invariances, the reduction in additionally required data can be substantial. Nevertheless, we show that our method yields the same optimal restoration function as supervised learning, which cannot be expected from purely self-supervised methods.
@@ 2.1 Problem formulation
The optimal restoration function $\hat{f}\colon\mathds{R}^{I}\to\mathds{R}^{I}$ is defined as the minimizer of the supervised loss function
$\mathcal{L}(f)\coloneqq\mathbb{E}\!\big[\|f(X_{\Omega})-X\|^{2}\big]\,$, (2.1)
where the minimum is taken over all equivalence classes (modulo almost-everywhere equality) of measurable functions $f\colon\mathds{R}^{I}\to\mathds{R}^{I}$. It is well known that the minimizer of (2.1) is given by the conditional expectation $\hat{f}=\mathbb{E}[X\mid X_{\Omega}]$. The conditional expectation is unique up to a set of $X_{\Omega}$-probability zero. Thus, $\hat{f}(z)$ is well defined whenever $z=x_{\Omega}=M_{\Omega}\odot x$ for some image $x\in\mathds{R}^{I}$ (again up to null sets). On the other hand, whenever $z$ does not vanish on $I\setminus\Omega$, the value $f(z)$ is not determined by (2.1) and can be chosen arbitrarily. For convenience, we still define $\hat{f}$ on the full space $\mathds{R}^{I}$, keeping in mind that (2.1) specifies it only on the range of the masking operator $M_{\Omega}$. This choice aligns with the notation commonly used in the self-supervision literature and also reflects that the neural network architectures $(f_{\theta})_{\theta\in\Theta}$ by default take inputs in $\mathds{R}^{I}$. The actual function we use for inference based on data $y$ in (1.1) is $\Phi_{\theta}(y)=(f_{\theta}\circ P_{\Omega}^{*})(y)=f_{\theta}(x_{\Omega})$; thus, it is not affected by how we extend $f_{\theta}$ outside the range of the masking operator $M_{\Omega}$.
Computing the exact minimizer in (2.1) requires exact knowledge of the joint distribution $\pi$, which is usually not available. In the supervised learning paradigm, the minimizer of (2.1) is instead approximated by minimizing the empirical loss $\sum_{n=1}^{N}\lVert f(P_{\Omega}^{*}(y_{n}))-x_{n}\rVert^{2}$ using samples $(y_{n},x_{n})\in\mathds{R}^{\Omega}\times\mathds{R}^{I}$ drawn independently from the joint distribution. This, however, requires fully sampled ground truth images $x_{n}$, which again might not be available in practice.
@@ 2.2 Self-supervised upsampling
In the context of image upsampling or inpainting, the self-supervision paradigm selects a second subsampling set $\Gamma\subseteq\Omega$, which further degrades the partially observed data. This procedure defines synthetic training pairs $(P_{\Lambda}y_{n},y_{n})$, where the original partially observed image $y_{n}$ serves as the ground truth, and $P_{\Lambda}y_{n}$ acts as the input to the model. In the context of image restoration with known masks, the main theoretical justification is that $\mathbb{E}[X\mid M_{\Omega}\odot X]$ can, in fact, be constructed from data $(P_{\Lambda}y,y)$.
For the reader’s convenience, we recall a main theoretical result in this context due to [9], where multiplicative variants of Noisier2Noise and their relation to SSDU (self-supervised learning via data undersampling, proposed in [16]) have been studied. A main recovery result states that
$M_{I\setminus(\Omega\cap\Lambda)}\odot\mathbb{E}[X\mid M_{\Lambda}\odot X_{\Omega}]\\ =M_{I\setminus(\Omega\cap\Lambda)}\odot\Bigl(\operatorname*{arg\,min}_{\Phi}\mathbb{E}\bigl[\lVert M_{I\setminus\Lambda}\odot\bigl(M_{\Omega}\odot\Phi(M_{\Lambda}\odot X_{\Omega})-X_{\Omega}\bigr)\rVert^{2}\mid M_{\Lambda}\odot X_{\Omega}\bigr]\Bigr)\,$, (2.2)
where $\Omega,\Lambda\subseteq I$ are assumed to be random subsets with $\mathbb{E}[M_{\Omega}]>0$ and $\mathbb{E}[M_{\Lambda}]<1$. (2.2) states that, on the pixels outside $\Omega\cap\Lambda$, the ideal estimator $\mathbb{E}[X\mid M_{\Lambda}\odot X_{\Omega}]$ for the given further masked data $M_{\Lambda}\odot X_{\Omega}$ can be found using only access to $X_{\Omega}$ instead of $X$ during training. In our opinion, this is quite remarkable. A key assumption is the randomness of the set $\Omega$, which, in expectation, covers the whole region $I$. In many practical scenarios, however, the set $\Omega$ may be deterministic and fixed and, moreover, may not cover the entire imaging domain. Our method, on the other hand, allows for a fixed, non-random, incomplete design. To obtain sufficient information, we exploit invariances in the distribution of $X$.
@@ 3.2 Main results
Let us briefly discuss Condition (3.9). Items (A1), (A2), and (A5) determine the setting we work in and the observed data, and are therefore not restrictive. By contrast, (A3) and (A4) may appear to impose potentially restrictive conditions on the supervision set $B$ and on the class $\mathcal{T}$. However, this is not the case; we consider the interplay between these conditions, as reflected in (A3) and (A4), to be quite natural. Our aim is to reduce the supervision set from the full set $I$ of pixels to a potentially small one, and the way to do so is to exploit invariances present in the dataset $X$. If there is no invariance, our method cannot reduce the supervision set. Formally, with only the identity mapping $T_{1}=\operatorname{Id}$ as an invariance, one must take $\Omega=I$; nothing is gained. If, however, there is at least one additional (nontrivial) invariance $T_{2}$, we may consider $\mathcal{T}=\{T_{1},T_{2}\}$. As a concrete example, let $T_{2}$ be a periodic shift by $N/2$ pixels in the horizontal dimension. Then we can take the supervision set $\Omega$ as the left half of the image, and (A3) and (A4) are satisfied. The more general case simply allows multiple invariances. While having more invariances may be harder to achieve in practice, in that case (A4) becomes easier to satisfy.
@@
The equivariance of $\mathbb{E}[X|X_{\Omega}]$ follows from Lemma 3.4. Further, from the $\mathcal{T}$-equivariance of $\mathbb{E}[X|X_{\Omega}]$ and conditions (A3), (A4), we get
@@
Recall that $\hat{f}=\mathbb{E}\!\left[X\mid X_{\Omega}\right]$ is the minimizer of the supervised loss (2.1), which is unique up to sets of $X_{\Omega}$-probability zero. Essential non-uniqueness can arise only when $\hat{f}$ is evaluated at elements $z$ that have zero probability of occurring as $X_{\Omega}$. We emphasize that in such a situation the right-hand side of (3.1) also does not provide additional information about $z$. This can be interpreted as follows: for strictly out-of-distribution samples, neither of the losses (2.1) or (3.1) provides any information and any prediction is feasible. Conversely, in all other cases where predictions are made on feasible elements, both loss functions yield the same unique minimizer.
@@ 3.3 Translation-invariant case
Let us discuss some practical implications and features of the equality (3.4). When we choose the supervision set $B\subsetneq I$, the optimal restoration function $\mathbb{E}\!\left[X\mid X_{\Omega}\right]$ can be learned using only instances of $X_{B}$, which form a smaller dataset than instances of the original $X$. In particular, this avoids collecting any data on the complementary domain $I\setminus B$, which is crucial if the complementary domain is inaccessible due to practical constraints, or if collecting data on different domains requires adjusting the device, which may be costly, time-consuming, and error-prone. Besides these practical advantages, it introduces the numerical challenge of minimizing over $\mathcal{T}$-equivariant functions $f$ rather than arbitrary functions. Several strategies are possible to guarantee $\mathcal{T}$-equivariance during numerical optimization. In the numerical implementation presented below, we start with a standard architecture (specifically, a U-Net) and then adjust the main building blocks (in this case, the convolutions) to be translation-equivariant, for example by enforcing appropriate boundary conditions. Other possible choices are currently under investigation and will be presented in future work.
@@ 4.2 Sparse-Dense OR-PAM
The partition condition (B4) is in this case readily verified. Choices without additional supervised data, such as $\tilde{B}=\Omega$ or $\tilde{B}=\Omega\cap B$, are not possible, as this would contradict (B4). Other invariances are also possible, for example, rotating the image by multiples of $90$ degrees; however, in this paper, we stick to translation invariances only. The numerical realization of these broader invariances presents additional challenges; appropriate strategies will be developed in future work.
Recall the fundamental difference between the observation set $\Omega$ and the supervision set $B$. The observation set $\Omega$ is required to be invariant under the transformations $(T_{\ell})_{\ell}$. The supervision set, on the other hand, is required to break this invariance: applying $(T_{\ell})_{\ell}$ to $B$ should cover the full image domain $I$. The aim of our sparse–dense sampling scheme is to reduce the size of the supervision region compared to the full supervision region $I$ used in standard supervised learning. To this end, we exploit the invariances $T_{1},T_{2},T_{3},T_{4}$.
@@ 4.3 Implementation details
To ensure that the network satisfies the condition of a $\mathcal{T}$-equivariant function, we analyze the translation-invariance properties of the architecture in detail and adapt elements when necessary. Specifically, this means the following.
  • Without boundary effects, convolutional layers are translation-equivariant because the same filter is applied with shared weights across all spatial positions, so a shift in the input produces an equivalent shift in the feature maps. When restricting to a finite domain with standard settings, this property holds only in the interior of the image; at the boundaries, padding and edge effects may break perfect equivariance. To overcome this issue, in our implementation all convolutional layers use padding_mode=circular, which ensures that the convolution operator is translation-equivariant with periodic boundary conditions. For convolutions with stride $s=1$, translation invariance holds for all integer shifts. However, the encoder path contains three strided convolutional layers (stride=2) acting as downsampling operators. Similarly, the decoder path contains three strided transposed convolutional layers (stride=2) acting as upsampling operators. For these layers, translation invariance is guaranteed only for translations that are multiples of the stride. Otherwise, aliasing effects occur, which break exact invariance. Since there are three downsampling operations in the encoder, the overall stride is $2\cdot 2\cdot 2=8$.
@@ 4.4 Numerical Studies
Figure 4.7: Visualization of superior performance of sparse-dense sampling over patch-wise supervised learning. From left to right: ground truth (image 1); downsampled image (image 2); upsampling with a patch-based supervised approach (image 3); upsampling with sparse-dense sampling (image 4).
@@ 5 Conclusion
Future work will focus on studying the influence of noise and exploring other invariances, such as rotation and mirroring. The experimental realization currently relies on translation equivariance, and extensions to rotations and other transforms would likely require different architectures and strategies. To investigate the boundaries of our method, we will also examine how far the sampling rate can be reduced; for example, recording only every third or fourth pixel in the observation and reducing the size of the supervision region. Finally, we plan to extend our method to random sampling and self-supervision with noisy data, where not all samples are available.